log base x (8/27)=3
Answer by stanbon(75887)

log base x (8/27)=3
-------
logx(8/27) = 3
-----
x^3 = (8/27)
---
x^3 = (2/3)^3
x = 2/3
=============
Cheers,
Stan H.
=============